# Physics-Informed Wireless Imaging with Implicit Neural Representation in RIS-Aided ISAC System

### v1

@@ 0 Algorithm GT INR INR FT CS
IV-B2 Imaging Result Comparison of INR with FT and CS Algorithms
We compare the imaging performance of FT [5], CS [13, 6], and proposed INR algorithms. To maintain consistent settings, only the TX-ROI-RIS-RX path is employed for imaging, TX and RX antenna numbers are both set to 1, and RIS phases are configured using a DFT codebook. The imaging distance is $D=40\lambda$. As shown in Table II, the proposed INR-based imager significantly outperforms traditional FT and CS methods. With only $K=500$ RIS phase variations, INR achieves high imaging quality. In contrast, even with $K=2500$ measurements, FT and CS generate degraded images due to strong pixel-wise channel correlations, which limit their ability to resolve fine details. This highlights the advantage of leveraging INR to learn intricate image features directly from CSI data. Furthermore, INR enables the generation of super-resolution images with pixel size $\xi_{\text{v}}$ much smaller than the resolution limit established in [5]. By comparison, FT and CS can only produce coarse approximations when $\xi_{\text{v}}=\lambda$. Therefore, the INR framework presents an effective alternative to conventional model-based imaging approaches, offering enhanced resolution and robustness in underdetermined sensing scenarios.
IV-B3 Imaging Performance versus Distance and RIS Phase Variation Number
@@
This subsection investigates the effects of distance $D$ and RIS phase variation number $K$ on imaging quality. The maximum training epoch is set to 1,000. The average MSE for 1,000 images using the proposed INR method are shown in Fig. 5. Several key insights can be drawn.
@@
First, the MSE results indicate that imaging quality initially improves and then degrades as the distance $D$ increases. This suggests the existence of an optimal imaging distance $D_{0}$, analogous to the focal distance in optical imaging, where performance is maximized. In general, $D_{0}$ decreases as $K$ increases, as illustrated by the dotted black line in Fig. 5. Therefore, collecting more CSI measurements allows the proposed INR-based imager to achieve optimal performance at shorter distances. These findings differ from traditional results [13, 5], which typically conclude that imaging performance deteriorates monotonically with increasing $D$.
@@
Second, when $D<D_{0}$, imaging performance is primarily constrained by the number of RIS phase variations $K$. In this region, the short imaging distance makes the ROI generate a large field of view to the RIS aperture, requesting a large $K$ to capture adequate information about the ROI, which resembles the scenario when a man observes a near and large object. As a result, a low $K$ leads to insufficient sensing perspectives and degraded performance, similar to the no-RIS scenario in Table I. Increasing $K$ from 5 to 80 can enrich the CSI measurements and ROI-related information, reducing MSE from around $10^{-1}$ to $10^{-3}$ at $D=20\lambda$.
@@
Third, when $D>D_{0}$, the influence of $K$ diminishes. For example, when $D$ approaches $200\lambda$, the MSE values for $K=5$ and $K=80$ become nearly identical. This indicates that the performance bottleneck shifts from RIS phase diversity to the increasing channel correlations between the RIS and ROI. These correlations make it more difficult to distinguish adjacent pixels as $D$ increases, consistent with analysis in [13, 5]. Consequently, imaging resolution deteriorates at larger distances. This condition also shares similarities with optical imaging systems where multiple observation may not significantly improve the resolution at large imaging distances.
V Conclusion
@@
This study presents a RIS-aided ISAC system in which the communication infrastructure is reused to enable target imaging. The INR technique is innovatively introduced for wireless imaging, where the NN architecture is specifically designed using positional encoding and a sine activation function to embed image information within the NN parameters. Under the supervision of physical models, DL techniques are employed to optimize the image representation, offering several advantages over traditional model-based and data-driven methods. Simulation results demonstrate that the proposed algorithm outperforms state-of-the-art methods by achieving high imaging accuracy with low measurement overhead.

### v2

@@ 0 Algorithm GT INR INR FT CS
We compare the imaging performance of FT [18], CS [17, 9, 16], and proposed INR algorithms. To maintain consistent simulation settings, only the TX-ROI-RIS-RX path is employed for image reconstruction, the numbers of antennas at the TX and RX are both set to 1, and the RIS phases are configured using a DFT codebook. The imaging distance is set to $D=40\lambda$. As shown in Table IV, the proposed INR-based imager significantly outperforms traditional FT and CS methods. With only $K=500$ RIS phase variations, INR achieves high imaging quality. In contrast, even with $K=2500$ measurements, FT and CS generate degraded images due to strong pixel-wise channel correlations, which limit their ability to resolve fine details. This highlights the advantage of leveraging DL techniques, as INR is capable of learning intricate image features directly from CSI data. Furthermore, INR enables the generation of super-resolution images with pixel size $\xi_{\text{v}}$ much smaller than the resolution limit established in [18]. By comparison, FT and CS can only produce coarse approximations when $\xi_{\text{v}}=\lambda$. Therefore, the INR framework presents an effective alternative to conventional model-based imaging approaches, offering enhanced resolution and robustness in underdetermined sensing scenarios.
VI-B4 Imaging Performance versus Distance and RIS Phase Variation Number
@@
Figure 10: SSIM versus imaging distance and RIS phase variation number.
@@
This subsection investigates the effects of imaging distance $D$ and the number of RIS phase variations $K$ on the imaging quality. To reduce simulation time, the maximum training epoch is set to 1,000. The average MSE and SSIM results for 1,000 images using the proposed INR algorithm are shown in Fig. 9 and Fig. 10, respectively. Several key insights can be drawn from these results.
@@
First, both MSE and SSIM indicate that imaging quality initially improves and then degrades as the distance $D$ increases from $20\lambda$ to $200\lambda$. This suggests the existence of an optimal imaging distance $D_{0}$, analogous to the focal distance in optical imaging, where performance is maximized. However, $D_{0}$ varies with $K$; for instance, $D_{0}=60\lambda$ when $K=5$, and $D_{0}=30\lambda$ when $K=40$. In general, $D_{0}$ decreases as $K$ increases, as illustrated by the dotted black line in Fig. 9. Therefore, collecting more CSI measurements allows the proposed INR-based imager to achieve optimal performance at shorter distances. These findings differ from traditional results [17, 18], which typically conclude that imaging performance deteriorates monotonically with increasing $D$.
@@
Second, when $D<D_{0}$, imaging performance is primarily constrained by the number of RIS phase variations $K$. In this region, the short imaging distance makes the ROI generate a large field of view to the RIS aperture, requesting a large $K$ to capture adequate information about the ROI, which resembles the scenario when a man observes a near and large object. As a result, a low $K$ leads to insufficient sensing perspectives and degraded performance, similar to the no-RIS scenario in Table II. Increasing $K$ from 5 to 80 can enrich CSI measurements and ROI-related information, reducing MSE from around $10^{-1}$ to $10^{-3}$ and increasing SSIM from 0.1 to 0.8 at $D=20\lambda$.
@@
Third, when $D>D_{0}$, the influence of $K$ diminishes. For example, when $D$ approaches $200\lambda$, the MSE and SSIM values for $K=5$ and $K=80$ become nearly identical. This indicates that the performance bottleneck shifts from RIS phase diversity to the increasing channel correlations between the RIS and ROI. These correlations make it more difficult to distinguish adjacent pixels as $D$ increases, consistent with analysis in [17, 18]. Consequently, imaging resolution deteriorates at larger distances. This condition also shares similarities with optical imaging systems where multiple observations may not significantly improve the resolution at large imaging distances.
VI-B5 Imaging Performance versus Noise Power and TX Antenna Number
Figure 11: MSE versus noise power and TX antenna number.
Figure 12: SSIM versus noise power and TX antenna number.
Fig. 11 and Fig. 12 illustrate the effects of varying noise power $P_{\text{n}}$ and the number of TX antennas $N_{\text{t}}$ on imaging performance, where $P_{\text{n}}$ may originates from channel estimation and background calibration errors. The number of RX antennas is set to $N_{\text{r}}=4$, and the transmit power is fixed at $P_{\text{t}}=20$ dBm. Results are averaged over 1,000 test images, using $K=40$ RIS phase variations and an imaging distance of $D=40\lambda$. The results show that the proposed algorithm is robust to noise when $P_{\text{n}}<-80$ dBm, with minimal impacts from additive noise. However, when $P_{\text{n}}>-80$ dBm, the imaging quality degrades as noise increases, leading to noise-dominated results as $P_{\text{n}}$ approaches 0 dBm. Under fixed transmit power, increasing the number of TX antennas improves performance in both MSE and SSIM metrics, despite the reduced per-antenna signal power. These improvements are attributed to the enhanced spatial resolution offered by the TX antenna array. Nevertheless, the performance gain diminishes as $N_{\text{t}}$ becomes large, indicating a trade-off between hardware cost and imaging quality.
VI-B6 Imaging Performance under TX Position Errors
The calculation of the physical model relies on the locations of the TX, RX, and RIS, which are typically measured manually and are thus susceptible to measurement errors. We have conducted simulations to evaluate the imaging performance against TX location accuracy ($\Delta$), and the results are shown in Table V. When $\Delta=10^{-2}\lambda$, the imaging result is comparable to that with no location errors. However, as $\Delta$ increases, the imaging results become progressively distorted. At $\Delta=10^{-0.5}\lambda$, significant noise artifacts appear, potentially obscuring the human body images. Consequently, high-precision antenna localization is essential for ensuring high-quality image generation. This vulnerability to location errors is also a common challenge for traditional imaging algorithms.
TABLE V: Imaging results with different TX position errors.
0 $\Delta$ ($\lambda$) 0 $10^{-2}$ $10^{-1.5}$ $10^{-1}$ $10^{-0.5}$
Result [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
VI-B7 INR-Based 3D Imaging
Figure 13: 3D imaging results: (a) Original 3D image; (b) Discretized GT 3D image for CSI generation; (c) Learned 3D image with coarse resolutions; (d) Rendered image with the resolution of $\lambda/5$; (e) Rendered image with the resolution of $\lambda/20$.
We further evaluate the 3D imaging ability of the proposed INR-based imaging method using a 3D human body model shown in Fig. 13(a). 3D imaging has significantly enlarged the size of the ROI, leading to a large image dimension $N_{\text{v}}$ and high computational complexity of the physical model. To accelerate image formulation, we employ a coarse grid for NN training with a voxel size of $\lambda/2$, discretizing the ROI of $1.6\times 0.4\times 2\ \text{m}^{3}$ into a point cloud image with $32\times 8\times 40$ voxels, as depicted in Fig. 13(b). Furthermore, the RIS size is set to $100\times 100$, and 10 subcarriers averagely spaced on a bandwidth of 400 MHz are employed. The training results are shown in Figs. 13(c-e), where Fig. 13(c) uses the same coarse grid for NN training, showing high consistency with Fig. 13(b) but low visualizability. Since the discretized training grid in Fig. 13(b) has lost substantial image details, the NN may not learn all the information involved in the original 3D model. However, the proposed INR-based imaging method can render images with arbitrary resolutions, and off-grid voxel values can be obtained by increasing the sampling grid density. Consequently, we can derive Fig. 13(d) and Fig. 13(e) with resolutions of $\lambda/5$ and $\lambda/20$, respectively. These images have presented smoother surfaces and more semantic details. Therefore, even trained with coarse grids, the proposed INR-based imaging method can still generate high-resolution images, which is significantly superior to traditional model-based and data-driven DL methods.
VI-B8 Imaging Performance under Multipath Interference
Figure 14: Performances of target imaging under multipath interference ${\mathbf{H}}_{\text{ima},k}$ in (24). “proposed” denotes the results obtained using the joint imaging and interference learning method in Sec. IV-A2, “compared” refers the results conducting imaging without learning ${\mathbf{H}}_{\text{ima},k}$, and “no interference” indicates the case with no multipath interference.
This subsection evaluates the effectiveness of the proposed imaging algorithm under multipath interference in Sec. IV-A. It is assumed that ROI-irrelevant multipaths have been repetitively measured and removed from the CSI measurements $\widehat{\mathbf{H}}_{\text{sen},k}$, following the background calibration approach in Sec. IV-A1. The ROI image is then reconstructed under the influence of the unknown multipath component $\mathbf{H}_{\text{roi-bg}}$ defined in (24). Based on (24) and (25), the signal-to-interference ratio (SIR) for imaging is defined as
$\text{SIR}=\frac{\|\mathbf{H}_{\text{tx-roi-rx}}+\mathbf{H}_{\text{tx-roi-ris-rx},k}+\mathbf{H}_{\text{tx-ris-roi-rx},k}\|_{2}^{2}}{\|\mathbf{H}_{\text{roi-bg}}\|_{2}^{2}}.$ (34)
Additionally, the interference recovery accuracy (IRA) is defined as
$\text{IRA}=\frac{\|\widehat{\mathbf{A}}\widehat{\bm{\sigma}}-\mathbf{A}\bm{\sigma}\|_{2}^{2}}{\|\mathbf{A}\bm{\sigma}\|_{2}^{2}},$ (35)
where $\widehat{\mathbf{A}}$ and $\widehat{\bm{\sigma}}$ denote the estimates of $\mathbf{A}$ and $\bm{\sigma}$, respectively. We consider $K=40$ and $D=40\lambda$, and compare the proposed joint imaging and interference learning method against conventional imaging without interference modeling.
As shown in Fig. 14, the proposed method significantly outperforms the baseline across various SIR levels. When the SIR exceeds 0 dB, the method can effectively optimize both the image representation $\mathcal{M}_{\bm{\theta}}$ and the multipath-related matrix $\mathbf{A}$, achieving imaging performance comparable to the interference-free case. However, performance gradually degrades as SIR falls below 0 dB. The matrix $\mathbf{A}$ becomes untrainable when SIR drops below -20 dB. These results indicate that the proposed INR-based imager remains effective under a partially known forward model, provided that the interference power is lower than that of the modeled components. In contrast, conventional imaging without interference learning is anticipated to perform well only when SIR exceeds 20 dB in the considered scenario, due to the mismatch between the assumed and actual multipath channels. Furthermore, the results labeled “compared” can also reveal the impacts of calibration errors in Sec. IV-A1, which are not learned by the proposed method.
VI-B9 Successive Imaging Performance for Dynamic Targets
This subsection evaluates the performance of successive imaging for dynamic targets using the prior-embedding method described in Sec. IV-B. We consider three successive time instants, during which two people move within the ROI. Their body shapes change over time, as shown in the “GT” row of Table VI. However, the three images remain highly correlated, allowing the image at each subsequent instant to benefit from the prior representation. At time instant $t=1$, we use 5,000 training epochs to obtain an accurate image with an SSIM of 0.8898. At the following two instants, only 200 training epochs are allowed in order to accelerate the imaging process and increase the frame rate. As shown in Table VI, training from scratch at later instants results in distorted images and low quality. In contrast, initializing the neural network with priors from time instant $t=1$ enables visually accurate reconstructions after just 200 epochs, with significantly higher PSNR and SSIM values compared to training from scratch.
Fig. 15 presents the test PSNR and SSIM values during training for the second time instant. The results show that training with priors can achieve the PSNR performance of training from scratch at epoch 2,000 and 5,000 by epoch 860 and 3,100, respectively. Similarly, the same SSIM achieved by training from scratch at epoch 2,000 (or 5,000) is reached with priors by epoch 60 (or 730). Therefore, embedding implicit priors from the initial instant significantly improves imaging quality with fewer training epochs, enabling efficient successive imaging for dynamic targets.
TABLE VI: Imaging results for dynamic targets.
0 Time instant $t$ 1 2 3
GT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Train from scratch Epoch 5,000 200 200
PSNR (dB) 24.0932 10.2783 10.3673
SSIM 0.8898 0.3653 0.3726
Result [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Train with priors Epoch / 200 200
PSNR (dB) / 18.4271 17.3824
SSIM / 0.8171 0.8278
Result / [Uncaptioned image] [Uncaptioned image]
Figure 15: Test PSNR and SSIM during the training process, comparing models trained from scratch and with priors.
Furthermore, we evaluate the prior-embedded performance gain across diverse image correlation levels. Specifically, we employ the image at $t=1$ in Table VI as the prior image. However, the subsequent target image is varied, and the SSIM metric is used to measure the image correlation between subsequent and prior images. The simulation results at epoch 600, presented in Fig. 16, reveal that the performance gain originating from the prior-embedded NN parameters gradually degrades as image correlation decreases. When the SSIM value between the prior and target images is reduced to 0.3, nearly the same PSNR performance is achieved regardless of whether prior information has been utilized. However, when the number of training epochs is lower than 600, the proposed prior-aided imaging method consistently exhibits a rapid convergence rate, enabling coarse image formulation within a short duration, as depicted in Fig. 15.
Figure 16: PSNR performance with/without image priors for different SSIM values.
VI-B10 Imaging-Augmented Communication Performance
We evaluate the enhancement of communication performance using high-quality imaging results obtained from the INR-based method. The transmit power is set to $P_{\text{t}}=0$ dBm, and we assume that the LOS path between TX and the user is not obstructed. The TX employs $N_{\text{t}}=2$ antennas. The RIS phase optimization algorithm proposed in [39] is adopted, and the RIS consists of $10\times 10$ elements to reduce the computational complexity. We respectively consider two scenarios where the user is either inside or outside the ROI.
When the user is located inside the ROI, it simultaneously serves as both the communication terminal and the target for sensing, as described in Sec. V-B. We assume that the rough user location is $40\lambda,0,0]^{\text{T}}$, which is used for RIS phase optimization without imaging results. However, the actual position of the communication device (the user hand) is $40\lambda,10\lambda,10\lambda]^{\text{T}}$, which can be identified using the INR-based high-quality image and computer vision techniques. Fig. 17(a) shows the SE under random and optimized RIS phase configurations. The results indicate that optimizing RIS phases based on inaccurate user positions may significantly degrade communication performance, potentially resulting in lower SE than with random RIS phase settings. In contrast, optimizing the RIS phases using the accurate terminal location obtained from the imaging results increases the SE by 2.76 bit/s/Hz compared to the random configuration.
(a) User inside the ROI
(b) User outside the ROI
Figure 17: Communication SE versus noise power. “case 1” denotes optimized RIS phases with information provided by imaging results, “case 2” denotes optimized RIS phases without information provided by imaging results, and “case 3” denotes random RIS phases.
When the user is located outside the ROI, the targets within the ROI provide additional multipaths to the communication channel, as explained in Sec. V-C. The user is assumed to be located at $30\lambda,-20\lambda,20\lambda]^{\text{T}}$, which has been previously estimated for RIS phase optimization. The corresponding SE results are illustrated in Fig. 17(b). It is observed that RIS phase optimization improves SE even when ROI-related multipaths are not considered in case 2, as their energy is typically low and may have limited influence on signal power. However, incorporating environmental information obtained from the imaging results in case 1 leads to further improvement in SE, thereby enhancing the communication experience. This performance gain is attainable only through the INR-based imager, which produces high-quality images, unlike traditional FT and CS methods that may not provide such information, as compared in Table IV.
VII Conclusion and Future Research Directions
@@
This study presents a RIS-aided ISAC system in which the communication infrastructure is reused to enable target imaging. The INR technique is introduced for wireless imaging, where the NN architecture is specifically designed using positional encoding and a sine activation function to embed image information within the NN parameters. Under the supervision of physical models, DL techniques are employed to optimize the image representation, offering several advantages over traditional model-based and data-driven methods. To improve imaging performance in practical environments, challenges such as multipath interference and dynamic target imaging are addressed. Furthermore, the imaging results are utilized to enhance communication SE. Simulation results demonstrate that the proposed algorithm outperforms state-of-the-art methods by achieving high imaging accuracy with low measurement overhead. Multipath interference can be effectively learned to support high-quality imaging, and the use of priors facilitates imaging for dynamic targets. Additionally, the communication performance of users located either inside or outside the ROI can be improved through the assistance of imaging results.
Our future work may include optimizing the NN with advanced positional encoding methods, activation functions, and accelerated training strategies. Moreover, the INR-based imaging algorithm is anticipated to be applied to diverse scenarios, including imaging using multiple observation views and multiple frequency bands. Validation through field trials conducted in live commercial systems is also required to verify the practical applicability of the proposed algorithms.
Appendix A
According to the channel models presented in Sec. II and Sec. V, we have $\mathbf{h}_{\text{com}}(\bm{\omega})=\mathbf{B}\bm{\omega}+\mathbf{c}$, where $\mathbf{B}=g_{\text{com}}g_{\text{ris}}\mathbf{H}_{\text{tx-ris}}\text{diag}(\mathbf{h}_{\text{ris-ue}})$ and $\mathbf{c}=\mathbf{h}_{\text{tx-ue}}$ when the user is inside the ROI, while $\mathbf{B}=g_{\text{com}}g_{\text{ris}}(\mathbf{H}_{\text{tx-ris}}\text{diag}(\mathbf{h}_{1})+\mathbf{H}_{1}\text{diag}(\mathbf{h}_{\text{ris-ue}}))$ and $\mathbf{c}=\mathbf{h}_{\text{tx-ue}}+\mathbf{h}_{\text{tx-roi-ue}}$ when the user is outside the ROI. Consequently, the problem (P5) can be rewritten as
(P6) $\max\limits_{\bm{\omega}}$ $\bm{\omega}^{\text{H}}\mathbf{B}^{\text{H}}\mathbf{B}\bm{\omega}+\bm{\omega}^{\text{H}}\mathbf{B}^{\text{H}}\mathbf{c}+\mathbf{c}^{\text{H}}\mathbf{B}\bm{\omega}$
s.t. $|\omega_{n_{\text{s}}}|=1,\ \ \forall n_{\text{s}}=1,\ldots,N_{\text{s}}.$
By introducing an auxiliary variable $z$, problem (P6) is equivalently written as
(P7) $\max\limits_{\bar{\bm{\omega}}}$ $\bar{\bm{\omega}}^{\text{H}}\mathbf{R}\bar{\bm{\omega}}$
s.t. $|\bar{\omega}_{n_{\text{s}}}|=1,\ \ \forall n_{\text{s}}=1,\ldots,N_{\text{s}}+1,$
where
$\mathbf{R}=\begin{bmatrix}\mathbf{B}^{\text{H}}\mathbf{B}&\mathbf{B}^{\text{H}}\mathbf{c}\\ \mathbf{c}^{\text{H}}\mathbf{B}&\mathbf{0}\end{bmatrix},\quad\bar{\bm{\omega}}=\begin{bmatrix}\bm{\omega}\\ z\end{bmatrix}$ (36)
Define $\mathbf{W}=\bar{\bm{\omega}}\bar{\bm{\omega}}^{\text{H}}$, we have $\bar{\bm{\omega}}^{\text{H}}\mathbf{R}\bar{\bm{\omega}}=\text{tr}(\mathbf{R}\bar{\bm{\omega}}\bar{\bm{\omega}}^{\text{H}})=\text{tr}(\mathbf{RW})$, where $\mathbf{W}\succeq 0$ and $\text{rank}(\mathbf{W})=1$ should be satisfied. Here, $\mathbf{W}\succeq 0$ means that the matrix $\mathbf{W}$ is positive semidefinite. By applying semidefinite relaxation, problem (P7) is reduced to
(P8) $\max\limits_{\mathbf{W}}$ $\text{tr}(\mathbf{RW})$
s.t. $\mathbf{W}_{n_{\text{s}},n_{\text{s}}}=1,\ \ \forall n_{\text{s}}=1,\ldots,N_{\text{s}}+1,$
$\mathbf{W}\succeq 0.$
Problem (P8) is a standard convex semidefinite program, and it can be optimally solved by convex optimization solvers such as CVX. Furthermore, additional steps are presented in [39] to construct a rank-one solution from the optimal higher-rank solution of problem (P8).